Together, Huey, Dewey and Louie earn 72, 000 a month. If Huey earns 6,000 more than Dewey and 9,000 more than Louie, how much does Huey earn?
Answer by hamsanash1981@gmail.com(151)


let huey earnings be x , then dewey earns = x-6000 and louie earns = x-9000
x+x-6000+x-9000=72000
3x -15000=72000
3x = 87000
x = 29000
huey earns 29000